# DCA: Graph-Guided Deep Embedding Clustering for Brain Atlases

### v1

@@ 15.4 Choice of gray matter mask
In the main text, the corresponding ROI masks are extracted from FreeSurfer’s aparc+aseg.mgz [49]. Here, we show the result (Table 18) by defining the ROI with the same mask to MMP [8]and Schaefer al [2]. DCA still shows higer homogeneity and silhouette coefficients compared to corresponding atlas.
@@ 17 Spatiotemporal masking strategy
To ensure that masked regions form contiguous blocks in the high-resolution volume, we first spatially downsample the original $96\times 96\times 96$ grid by a factor of 16 along each axis, yielding a coarse $6\times 6\times 6$ volume (time dimension $T$ unchanged). We then sample two independent binary masks on this downsampled grid: a spatial mask with fraction $x_{s}$ of voxels set to zero, and a temporal mask with fraction $x_{t}$ of frames set to zero. We choose $x_{s}$ and $x_{t}$ so that the overall masking ratio satisfies
@@
where $r$ is the desired fraction of masked spatiotemporal volume (e.g. $r=0.8$ for 80% masking [57]). Finally, we upsample these binary masks back to the original resolution by expanding each downsampled voxel mask to a $16^{3}$ block in space and each temporal mask entry to the corresponding contiguous frames. Applying the resulting mask to the full-resolution data yields large, continuous spatiotemporal occlusions, encouraging the encoder to reconstruct missing patches using both local and long-range context.

### v2

@@ 15.4 Choice of gray matter mask
In the main text, the corresponding ROI masks are extracted from FreeSurfer’s aparc+aseg.mgz [48]. Here, we show the result (Table 18) by defining the ROI with the same mask to MMP [8]and Schaefer et al [2]. DCA still shows higher homogeneity and silhouette coefficients compared to corresponding atlas.
@@ 17 Spatiotemporal masking strategy
To ensure that masked regions form contiguous blocks in the high‐resolution volume, we first spatially downsample the original $96\times 96\times 96$ grid by a factor of 16 along each axis, yielding a coarse $6\times 6\times 6$ volume (time dimension $T$ unchanged). We then sample two independent binary masks on this downsampled grid: a spatial mask with fraction $x_{s}$ of voxels set to zero, and a temporal mask with fraction $x_{t}$ of frames set to zero. We choose $x_{s}$ and $x_{t}$ so that the overall masking ratio satisfies
@@
where $r$ is the desired fraction of masked spatiotemporal volume (e.g. $r=0.8$ for 80% masking [57]). Finally, we upsample these binary masks back to the original resolution by expanding each downsampled voxel mask to a $16^{3}$ block in space and each temporal mask entry to the corresponding contiguous frames. Applying the resulting mask to the full‐resolution data yields large, continuous spatiotemporal occlusions, encouraging the encoder to reconstruct missing patches using both local and long‐range context.
@@ 18 Train and validation loss
19 Data organization
There are two parts in our repository, DCA and ATlaScore. Please follow https://github.com/ncclab-sustech/DCA for details.
19.1 DCA
The codes for pretraining and personalized clustering. One example subject is provided.
DCA/
    data/
    fmri/
    mask/
    sub_test.txt
    data_preparation.ipynb
    results/
    demo/
    ablation_fmri.py
    main.py
    utils.py
    swin_unetr.py
19.2 AtlaScore
The codes for evaluating atlases. The volumetric Homogeneity and Silhouette evaluations, as well as the surface-based DCBC evaluation, are provided. All 12 downstream tasks are implemented in downstream/, with FC features for DCA100 provided as an example.
AtlaScore/
    similarity/
    compute_adj.py
    eva.py
    eva_DCBC.py
    downstream/
    docs/
    fc_data/
    nii_data/
    downstream.py
    demo.ipynb